# Measuring Vision-Language STEM Skills of Neural Models

### v1

@@ Abstract
We introduce a new challenge to test the STEM skills of neural models. The problems in the real world often require solutions, combining knowledge from STEM (science, technology, engineering, and math). Unlike existing datasets, our dataset requires the understanding of multimodal vision-language information of STEM. Our dataset features one of the largest and most comprehensive datasets for the challenge. It includes $448$ skills and $1,073,146$ questions spanning all STEM subjects. Compared to existing datasets that often focus on examining expert-level ability, our dataset includes fundamental skills and questions designed based on the K-12 curriculum. We also add state-of-the-art foundation models such as CLIP and GPT-3.5-Turbo to our benchmark. Results show that the recent model advances only help master a very limited number of lower grade-level skills ($2.5\%$ in the third grade) in our dataset. In fact, these models are still well below (averaging $54.7\%$) the performance of elementary students, not to mention near expert-level performance. To understand and increase the performance on our dataset, we teach the models on a training split of our dataset. Even though we observe improved performance, the model performance remains relatively low compared to average elementary students. To solve STEM problems, we will need novel algorithmic innovations from the community.
@@ A.1 Analysis
Counting
Count to 10, Count shapes in rows, Count sides and corners …
Geometry
Classify triangles, Identify symmetry, Identify shapes …
Time
Match times, Identify A.M./P.M., Read a calendar …
@@
(a) Number of skills.
(b) Skill comparison between STEM and IconQA.
@@ B.2 Detailed Experimental Analysis
Figure 18: Results on questions with different lengths.
@@ B.3 Error Analysis
Description: The school district compared how many swings each elementary school has. Which school has the fewest swings? Picture: [Uncaptioned image] Choices: [Shoreline Elementary, Hillside Elementary, Valley Elementary, Lincoln Elementary, ] Answer index: 2 Prediction: 0
@@
Description: Paul kept a log of how many minutes he spent practicing ice skating over the past 4 days. On which day did Paul practice the least? Picture: [Uncaptioned image] Choices: [Tuesday, Wednesday, Thursday, Friday, ] Answer index: 3 Prediction: 2
@@
Description: Look at the following schedule: Which meeting ends at 12:00 P.M.? Picture: [Uncaptioned image] Choices: [the city council meeting, the construction permit meeting, the parking meter meeting, the police meeting, ] Answer index: 0 Prediction: 2
@@
Description: What fraction of a turn is this angle? Picture: [Uncaptioned image] Choices: [3/4, 1 full turn, 1/2, 1/4, ] Answer index: 2 Prediction: 3
@@
Description: What is this? Picture: [Uncaptioned image] Choices: [a line segment, a ray, a line, a point, ] Answer index: 1 Prediction: 0
@@
Description: This passage describes the antenna type trait in fruit flies: Most fruit flies have a pair of antennae on their head. But, some flies appear to have an extra pair of legs on their head instead! These flies have a mutation, or change, in a gene that affects body development. This mutation makes the cells in the fly’s head form mutated antennae that are like legs. In a group of fruit flies, some individuals have mutated antennae and others have normal antennae. In this group, the gene for the antenna type trait has two alleles. The allele for normal antennae (a) is recessive to the allele for mutated antennae (A). This Punnett square shows a cross between two fruit flies. What is the expected ratio of offspring with normal antennae to offspring with mutated antennae? Choose the most likely ratio. Picture: [Uncaptioned image] Choices: [0:4, 3:1, 2:2, 1:3, 4:0, ] Answer index: 0 Prediction: 3
@@
Description: In a group of tomato plants, some individuals have smooth fruit and others have fuzzy fruit. In this group, the gene for the fruit texture trait has two alleles. The allele for smooth fruit (F) is dominant over the allele for fuzzy fruit (f). This Punnett square shows a cross between two tomato plants. What is the probability that a tomato plant produced by this cross will be homozygous recessive for the fruit texture gene? Picture: [Uncaptioned image] Choices: [0/4, 1/4, 2/4, 3/4, 4/4, ] Answer index: 0 Prediction: 3
@@
Description: Two identical blocks are heated to different temperatures. The blocks are placed so that they touch, and heat begins to flow between the blocks. The pair of blocks is insulated, so no energy escapes. Later, the temperature of each block is measured again. Which pair of temperatures is possible? Picture: [Uncaptioned image] Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1 Prediction: 0
@@
Description: Two magnets are placed as shown. Hint: Magnets that attract pull together. Magnets that repel push apart. Picture: [Uncaptioned image] Choices: [attract, repel, ] Answer index: 1 Prediction: 0
@@
Description: Two solid blocks are at different temperatures. The blocks are touching. Which picture shows how heat will move? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0 Prediction: 1
@@ Appendix C Additional Related Work
Figure 25: Accuracy per skill on math (part 3).
Figure 26: Accuracy per skill on science.
Figure 27: Accuracy per skill on technology.
@@ Appendix D Summary of Skills
Subject: Technology Description: This is a(n old) logo of which famous app or program? Picture: [Uncaptioned image] Choices: [Microsoft Office Outlook, Microsoft Office OneDrive, OfficeSuite Pro, Opera, ] Answer index: 3
Subject: Technology Description: What kind of computer component do you see here? Picture: [Uncaptioned image] Choices: [TV Tuner Card, PC Card, Motherboard, Modem Card, ] Answer index: 2
Subject: Technology Description: This is (part of) a (former) logo of which computer related brand? Picture: [Uncaptioned image] Choices: [ASRock, Amiga Inc., Arctic, ATI Technologies, ] Answer index: 3
Subject: Technology Description: This is (part of) a (former) logo of which computer related brand? Picture: [Uncaptioned image] Choices: [Fujitsu, Samsung, Iiyama, Brother, ] Answer index: 2
@@
Subject: Technology Description: This is (part of) a (former) logo of which computer related brand? Picture: [Uncaptioned image] Choices: [Xiaomi, Cisco, Intel, Wii, ] Answer index: 3
Subject: Technology Description: What kind of computer component do you see here? Picture: [Uncaptioned image] Choices: [Display Adapter/Video Card, PC Card, Power Supply Unit, Hard Disk Drive, ] Answer index: 3
Subject: Technology Description: What meaning or function is usually associated with this web interface symbol? Picture: [Uncaptioned image] Choices: [Paste, Search, Tip/Idea, Calendar/Event, ] Answer index: 2
Subject: Technology Description: This is a(n old) logo of which famous app or program? Picture: [Uncaptioned image] Choices: [YouTube Music, Beats Music, MX Player, YouTube, ] Answer index: 2
Subject: Technology Description: What kind of computer related plug or port do you see here? Picture: [Uncaptioned image] Choices: [USB type-C plug, DVI plug (type D), HDMI plug, 3.5mm Audio Cable plug, ] Answer index: 0
Subject: Technology Description: Identify this font type Picture: [Uncaptioned image] Choices: [Lucida MT, News Gothic MT, Fixedsys, Courier New, ] Answer index: 2
Subject: Technology Description: Identify this font type Picture: [Uncaptioned image] Choices: [Commercial Script BT, Brush Script MT, Vivaldi D, ShelleyVolante BT, ] Answer index: 3
Subject: Technology Description: Identify this font type Picture: [Uncaptioned image] Choices: [Garamond, Times New Roman, Courier New, Georgia, ] Answer index: 3
@@
Subject: Technology Description: This is (part of) a (former) logo of which computer related brand? Picture: [Uncaptioned image] Choices: [BenQ, Lexmark, Creative Technology, Lenovo, ] Answer index: 2
Subject: Technology Description: Identify this font type Picture: [Uncaptioned image] Choices: [Webdings, Courier, Impact, System, ] Answer index: 3
Subject: Technology Description: Identify this font type Picture: [Uncaptioned image] Choices: [Serifa BT, Stylus ITC, Calisto MT, Tempus Sans ITC, ] Answer index: 0
Subject: Technology Description: What meaning or function is usually associated with this web interface symbol? Picture: [Uncaptioned image] Choices: [Pin/Make something sticky, Storage for deleted files, Options/Settings, Print (preview), ] Answer index: 1
Subject: Technology Description: What meaning or function is usually associated with this web interface symbol? Picture: [Uncaptioned image] Choices: [Zoom in, Help, Like something, Link select, ] Answer index: 3
Subject: Technology Description: What meaning or function is usually associated with this web interface symbol? Picture: [Uncaptioned image] Choices: [Apply, Options/Settings, Reload/Refresh, Download, ] Answer index: 2
Subject: Technology Description: What type of video game console do you see here? Picture: [Uncaptioned image] Choices: [Mattel Intellivision, Sega Master System, Magnavox Odyssey 2, Atari 5200, ] Answer index: 3
@@
Subject: Technology Description: What meaning or function is usually associated with this web interface symbol? Picture: [Uncaptioned image] Choices: [Find, Delete, Attachment, Calendar/Event, ] Answer index: 0
@@
Subject: Engineer Description: Select the gloves. Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 0
@@
Subject: Engineer Description: iption: In this experiment, which were part of an experimental group? The passage below describes an experiment. Lucy and Erik were taking a snowboarding class. During the class, their instructor said they would go faster if they applied wax to the undersides of their snowboards. After the class, Lucy applied a thin layer of wax to the underside of a snowboard and rode the board straight down a hill. Then, she removed the wax and rode the snowboard straight down the hill again. Erik timed how long each ride took. Lucy repeated these rides on four other snowboards, alternating whether she first rode with or without wax. Picture: [Uncaptioned image] Choices: [the snowboards with wax removed, the snowboards with wax added, ] Answer index: 1
Subject: Engineer Description: Select the test tube. Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 2
Subject: Engineer Description: Select the funnel. Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 2
Subject: Engineer Description: Select the round-bottom flask. Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 1
@@
Subject: Engineer Description: iption: In this experiment, which were part of an experimental group? The passage below describes an experiment. Kimberly grew roses for a flower shop. One day, she noticed tumor-like growths on her rose stems. She could tell that the plants had crown gall disease, which is caused by a type of bacteria. She knew that allicin, a chemical in garlic, can kill bacteria. Kimberly wondered if spraying her plants with garlic juice would prevent more tumors from forming on her plants. Once a day, Kimberly sprayed garlic juice on ten infected plants and left another 10 infected plants unsprayed. After one month, she compared the number of new tumors on plants in the two groups. Picture: [Uncaptioned image] Choices: [the roses sprayed with garlic juice, the roses that were not sprayed, ] Answer index: 0
@@
Subject: Engineer Description: iption: Which of the following could Kendra’s test show? Wind turbines use wind power to produce electricity. Kendra was a materials engineer who designed wind turbines. She wanted to design a new turbine that would produce 10% more electricity than older wind turbines. She thought that a turbine made from lightweight material would turn more easily and produce more electricity. So, Kendra created a computer model of a turbine made from lightweight material. Then she used the model to calculate how much more electricity the new turbine could produce compared to the older turbines. The passage below describes how the engineering-design process was used to test a solution to a problem. Read the passage. Then answer the question below. Picture: [Uncaptioned image] Choices: [how much the new turbine would weigh, whether the new turbine could produce 10% more electricity, if the new turbine could turn easily, ] Answer index: 1
@@
Subject: Engineer Description: iption: In this experiment, which were part of an experimental group? The passage below describes an experiment. Isaac and his friend Belle flew nylon kites on the beach. They wondered if putting a tail on a kite would affect how well the kite flew. Isaac flew a kite that did not have a tail for five minutes. Then, he attached a four-foot-long tail and flew the kite for five more minutes. Isaac repeated this with three similar kites, alternating whether he started the kite with or without a tail. During each flight, Belle counted the number of times the kite crashed to the ground. Picture: [Uncaptioned image] Choices: [the kites without tails, the kites with tails, ] Answer index: 1
@@
Subject: Engineer Description: iption: Identify the question that Bryant and Lamar’s experiment can best answer. The passage below describes an experiment. Read the passage and then follow the instructions below. Bryant placed a ping pong ball in a catapult, pulled the catapult’s arm back to a 45° angle, and launched the ball. Then, Bryant launched another ping pong ball, this time pulling the catapult’s arm back to a 30° angle. With each launch, his friend Lamar measured the distance between the catapult and the place where the ball hit the ground. Bryant and Lamar repeated the launches with ping pong balls in four more identical catapults. They compared the distances the balls traveled when launched from a 45° angle to the distances the balls traveled when launched from a 30° angle. Picture: [Uncaptioned image] Choices: [Do ping pong balls stop rolling along the ground sooner after being launched from a 30° angle or a 45° angle?, Do ping pong balls travel farther when launched from a 30° angle compared to a 45° angle?, ] Answer index: 1
Subject: Engineer Description: Select the Erlenmeyer flask. Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 3
@@
Subject: Engineer Description: iption: Which of the following could Ivan’s test show? Ivan was a landscape architect who was hired to design a new city park. The city council wanted the park to have space for outdoor concerts and to have at least 20% of the park shaded by trees. Ivan thought the concert area should be at least 150 meters from the road so traffic noise didn’t interrupt the music. He developed three possible designs for the park with the concert area in a different location in each design. Then, he tested each design by measuring the distance between the road and the concert area. The passage below describes how the engineering-design process was used to test a solution to a problem. Read the passage. Then answer the question below. Picture: [Uncaptioned image] Choices: [if at least 20% of the park would be shaded by trees in each design, which design would have the greatest distance between the concert area and the road, which design would have the least traffic noise in the concert area, ] Answer index: 1
Subject: Engineer Description: Select the beaker. Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 3
@@
Subject: Engineer Description: iption: Identify the question that Zeke’s experiment can best answer. The passage below describes an experiment. Read the passage and then follow the instructions below. Zeke divided 40 unripe bananas evenly among eight paper bags and sealed the bags. He poked 20 small holes in four of the bags and left the other four without holes. He kept the bags at room temperature for three days. Then, Zeke opened the bags and counted the number of brown spots on each banana. He compared the average number of brown spots on bananas from bags with holes to the average number of brown spots on bananas from bags without holes. Picture: [Uncaptioned image] Choices: [Do bananas develop more brown spots if they are kept in bags with holes compared to bags without holes?, Do bananas develop more brown spots when they are kept at room temperature compared to in a cold refrigerator?, ] Answer index: 0
@@
Subject: Engineer Description: iption: Hint: An independent variable is a variable whose effect you are investigating. A dependent variable is a variable that you measure. Which of the following was an independent variable in this experiment? The passage below describes an experiment. Read the passage and think about the variables that are described. Tyler designed an electric circuit to test how well different types of metal conduct electricity. The circuit included a battery, a light bulb, wires, and clips that could be attached to a sheet of metal. If the metal conducted electricity poorly, the light bulb would appear dim. If the metal conducted electricity well, the light bulb would appear bright. Tyler collected nine equally sized sheets of metal: three sheets of copper, three sheets of iron, and three sheets of aluminum. He used the clips to attach each metal sheet, one sheet at a time, to the circuit. For each sheet, Tyler used a light meter to measure how much light the bulb produced. Picture: [Uncaptioned image] Choices: [the amount of light produced by the light bulb, the type of metal sheet used in the circuit, ] Answer index: 1
@@
Subject: Engineer Description: iption: Identify the question that Devon’s experiment can best answer. The passage below describes an experiment. Read the passage and then follow the instructions below. Devon poured four ounces of water into each of six glasses. Devon dissolved one tablespoon of salt in each of three glasses, and did not add salt to the other three. Then, Devon placed an egg in one glass and observed if the egg floated. She removed the egg and dried it. She repeated the process with the other five glasses, recording each time if the egg floated. Devon repeated this test with two more eggs and counted the number of times the eggs floated in fresh water compared to salty water. Picture: [Uncaptioned image] Choices: [Does the amount of water in a glass affect whether eggs sink or float in the water?, Are eggs more likely to float in fresh water or salty water?, ] Answer index: 1

### v2

@@ Abstract
We introduce a new challenge to test the STEM skills of neural models. The problems in the real world often require solutions, combining knowledge from STEM (science, technology, engineering, and math). Unlike existing datasets, our dataset requires the understanding of multimodal vision-language information of STEM. Our dataset features one of the largest and most comprehensive datasets for the challenge. It includes $448$ skills and $1,073,146$ questions spanning all STEM subjects. Compared to existing datasets that often focus on examining expert-level ability, our dataset includes fundamental skills and questions designed based on the K-12 curriculum. We also add state-of-the-art foundation models such as CLIP and GPT-3.5-Turbo to our benchmark. Results show that the recent model advances only help master a very limited number of lower grade-level skills ($2.5\%$ in the third grade) in our dataset. In fact, these models are still well below (averaging $54.7\%$) the performance of elementary students, not to mention near expert-level performance. To understand and increase the performance on our dataset, we teach the models on a training split of our dataset. Even though we observe improved performance, the model performance remains relatively low compared to average elementary students. To solve STEM problems, we will need novel algorithmic innovations from the community.  [^1] The dataset and leaderboard are available at https://huggingface.co/datasets/stemdataset/STEM.
@@ A.1 Analysis
Counting Count to 10, Count shapes in rows, Count sides and corners …
Geometry Classify triangles, Identify symmetry, Identify shapes …
Time Match times, Identify A.M./P.M., Read a calendar …
@@ B.3 Error Analysis
Description: The school district compared how many swings each elementary school has. Which school has the fewest swings?
Picture: [Uncaptioned image]
Choices: [Shoreline Elementary, Hillside Elementary, Valley Elementary, Lincoln Elementary, ]
Answer index: 2
Prediction: 0
@@
Description: Paul kept a log of how many minutes he spent practicing ice skating over the past 4 days. On which day did Paul practice the least?
Picture: [Uncaptioned image]
Choices: [Tuesday, Wednesday, Thursday, Friday, ]
Answer index: 3
Prediction: 2
@@
Description: Look at the following schedule: Which meeting ends at 12:00 P.M.?
Picture: [Uncaptioned image]
Choices: [the city council meeting, the construction permit meeting, the parking meter meeting, the police meeting, ]
Answer index: 0
Prediction: 2
@@
Description: What fraction of a turn is this angle?
Picture: [Uncaptioned image]
Choices: [3/4, 1 full turn, 1/2, 1/4, ]
Answer index: 2
Prediction: 3
@@
Description: What is this?
Picture: [Uncaptioned image]
Choices: [a line segment, a ray, a line, a point, ]
Answer index: 1
Prediction: 0
@@
Description: This passage describes the antenna type trait in fruit flies: Most fruit flies have a pair of antennae on their head. But, some flies appear to have an extra pair of legs on their head instead! These flies have a mutation, or change, in a gene that affects body development. This mutation makes the cells in the fly’s head form mutated antennae that are like legs. In a group of fruit flies, some individuals have mutated antennae and others have normal antennae. In this group, the gene for the antenna type trait has two alleles. The allele for normal antennae (a) is recessive to the allele for mutated antennae (A). This Punnett square shows a cross between two fruit flies. What is the expected ratio of offspring with normal antennae to offspring with mutated antennae? Choose the most likely ratio.
Picture: [Uncaptioned image]
Choices: [0:4, 3:1, 2:2, 1:3, 4:0, ]
Answer index: 0
Prediction: 3
@@
Description: In a group of tomato plants, some individuals have smooth fruit and others have fuzzy fruit. In this group, the gene for the fruit texture trait has two alleles. The allele for smooth fruit (F) is dominant over the allele for fuzzy fruit (f). This Punnett square shows a cross between two tomato plants. What is the probability that a tomato plant produced by this cross will be homozygous recessive for the fruit texture gene?
Picture: [Uncaptioned image]
Choices: [0/4, 1/4, 2/4, 3/4, 4/4, ]
Answer index: 0
Prediction: 3
@@
Description: Two identical blocks are heated to different temperatures. The blocks are placed so that they touch, and heat begins to flow between the blocks. The pair of blocks is insulated, so no energy escapes. Later, the temperature of each block is measured again. Which pair of temperatures is possible?
Picture: [Uncaptioned image]
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Prediction: 0
@@
Description: Two magnets are placed as shown. Hint: Magnets that attract pull together. Magnets that repel push apart.
Picture: [Uncaptioned image]
Choices: [attract, repel, ]
Answer index: 1
Prediction: 0
@@
Description: Two solid blocks are at different temperatures. The blocks are touching. Which picture shows how heat will move?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Prediction: 1
@@ Appendix D Summary of Skills
Subject: Technology
Description: This is a(n old) logo of which famous app or program?
Picture: [Uncaptioned image]
Choices: [Microsoft Office Outlook, Microsoft Office OneDrive, OfficeSuite Pro, Opera, ]
Answer index: 3
Subject: Technology
Description: What kind of computer component do you see here?
Picture: [Uncaptioned image]
Choices: [TV Tuner Card, PC Card, Motherboard, Modem Card, ]
Answer index: 2
Subject: Technology
@@
Description: This is (part of) a (former) logo of which computer related brand?
Picture: [Uncaptioned image]
Choices: [ASRock, Amiga Inc., Arctic, ATI Technologies, ]
Answer index: 3
Subject: Technology
@@
Description: This is (part of) a (former) logo of which computer related brand?
Picture: [Uncaptioned image]
Choices: [Fujitsu, Samsung, Iiyama, Brother, ]
Answer index: 2
Subject: Technology
Description: This is (part of) a (former) logo of which computer related brand?
Picture: [Uncaptioned image]
Choices: [Xiaomi, Cisco, Intel, Wii, ]
Answer index: 3
Subject: Technology
Description: What kind of computer component do you see here?
Picture: [Uncaptioned image]
Choices: [Display Adapter/Video Card, PC Card, Power Supply Unit, Hard Disk Drive, ]
Answer index: 3
Subject: Technology
@@
Description: What meaning or function is usually associated with this web interface symbol?
Picture: [Uncaptioned image]
Choices: [Paste, Search, Tip/Idea, Calendar/Event, ]
Answer index: 2
Subject: Technology
Description: This is a(n old) logo of which famous app or program?
Picture: [Uncaptioned image]
Choices: [YouTube Music, Beats Music, MX Player, YouTube, ]
Answer index: 2
Subject: Technology
Description: What kind of computer related plug or port do you see here?
Picture: [Uncaptioned image]
Choices: [USB type-C plug, DVI plug (type D), HDMI plug, 3.5mm Audio Cable plug, ]
Answer index: 0
Subject: Technology
Description: Identify this font type
Picture: [Uncaptioned image]
Choices: [Lucida MT, News Gothic MT, Fixedsys, Courier New, ]
Answer index: 2
Subject: Technology
Description: Identify this font type
Picture: [Uncaptioned image]
Choices: [Commercial Script BT, Brush Script MT, Vivaldi D, ShelleyVolante BT, ]
Answer index: 3
Subject: Technology
Description: Identify this font type
Picture: [Uncaptioned image]
Choices: [Garamond, Times New Roman, Courier New, Georgia, ]
Answer index: 3
Subject: Technology
Description: This is (part of) a (former) logo of which computer related brand?
Picture: [Uncaptioned image]
Choices: [BenQ, Lexmark, Creative Technology, Lenovo, ]
Answer index: 2
Subject: Technology
Description: Identify this font type
Picture: [Uncaptioned image]
Choices: [Webdings, Courier, Impact, System, ]
Answer index: 3
Subject: Technology
Description: Identify this font type
Picture: [Uncaptioned image]
Choices: [Serifa BT, Stylus ITC, Calisto MT, Tempus Sans ITC, ]
Answer index: 0
Subject: Technology
Description: What meaning or function is usually associated with this web interface symbol?
Picture: [Uncaptioned image]
Choices: [Pin/Make something sticky, Storage for deleted files, Options/Settings, Print (preview), ]
Answer index: 1
Subject: Technology
Description: What meaning or function is usually associated with this web interface symbol?
Picture: [Uncaptioned image]
Choices: [Zoom in, Help, Like something, Link select, ]
Answer index: 3
Subject: Technology
Description: What meaning or function is usually associated with this web interface symbol?
Picture: [Uncaptioned image]
Choices: [Apply, Options/Settings, Reload/Refresh, Download, ]
Answer index: 2
Subject: Technology
Description: What type of video game console do you see here?
Picture: [Uncaptioned image]
Choices: [Mattel Intellivision, Sega Master System, Magnavox Odyssey 2, Atari 5200, ]
Answer index: 3
Subject: Technology
Description: What meaning or function is usually associated with this web interface symbol?
Picture: [Uncaptioned image]
Choices: [Find, Delete, Attachment, Calendar/Event, ]
Answer index: 0
@@
Subject: Engineer
Description: Select the gloves.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Engineer
@@
Description: iption: In this experiment, which were part of an experimental group? The passage below describes an experiment. Lucy and Erik were taking a snowboarding class. During the class, their instructor said they would go faster if they applied wax to the undersides of their snowboards. After the class, Lucy applied a thin layer of wax to the underside of a snowboard and rode the board straight down a hill. Then, she removed the wax and rode the snowboard straight down the hill again. Erik timed how long each ride took. Lucy repeated these rides on four other snowboards, alternating whether she first rode with or without wax.
Picture: [Uncaptioned image]
Choices: [the snowboards with wax removed, the snowboards with wax added, ]
Answer index: 1
Subject: Engineer
Description: Select the test tube.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 2
Subject: Engineer
Description: Select the funnel.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 2
Subject: Engineer
Description: Select the round-bottom flask.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Engineer
@@
Description: iption: In this experiment, which were part of an experimental group? The passage below describes an experiment. Kimberly grew roses for a flower shop. One day, she noticed tumor-like growths on her rose stems. She could tell that the plants had crown gall disease, which is caused by a type of bacteria. She knew that allicin, a chemical in garlic, can kill bacteria. Kimberly wondered if spraying her plants with garlic juice would prevent more tumors from forming on her plants. Once a day, Kimberly sprayed garlic juice on ten infected plants and left another 10 infected plants unsprayed. After one month, she compared the number of new tumors on plants in the two groups.
Picture: [Uncaptioned image]
Choices: [the roses sprayed with garlic juice, the roses that were not sprayed, ]
Answer index: 0
Subject: Engineer
@@
Description: iption: Which of the following could Kendra’s test show? Wind turbines use wind power to produce electricity. Kendra was a materials engineer who designed wind turbines. She wanted to design a new turbine that would produce 10% more electricity than older wind turbines. She thought that a turbine made from lightweight material would turn more easily and produce more electricity. So, Kendra created a computer model of a turbine made from lightweight material. Then she used the model to calculate how much more electricity the new turbine could produce compared to the older turbines. The passage below describes how the engineering-design process was used to test a solution to a problem. Read the passage. Then answer the question below.
Picture: [Uncaptioned image]
Choices: [how much the new turbine would weigh, whether the new turbine could produce 10% more electricity, if the new turbine could turn easily, ]
Answer index: 1
Subject: Engineer
@@
Description: iption: In this experiment, which were part of an experimental group? The passage below describes an experiment. Isaac and his friend Belle flew nylon kites on the beach. They wondered if putting a tail on a kite would affect how well the kite flew. Isaac flew a kite that did not have a tail for five minutes. Then, he attached a four-foot-long tail and flew the kite for five more minutes. Isaac repeated this with three similar kites, alternating whether he started the kite with or without a tail. During each flight, Belle counted the number of times the kite crashed to the ground.
Picture: [Uncaptioned image]
Choices: [the kites without tails, the kites with tails, ]
Answer index: 1
Subject: Engineer
@@
Description: iption: Identify the question that Bryant and Lamar’s experiment can best answer. The passage below describes an experiment. Read the passage and then follow the instructions below. Bryant placed a ping pong ball in a catapult, pulled the catapult’s arm back to a 45° angle, and launched the ball. Then, Bryant launched another ping pong ball, this time pulling the catapult’s arm back to a 30° angle. With each launch, his friend Lamar measured the distance between the catapult and the place where the ball hit the ground. Bryant and Lamar repeated the launches with ping pong balls in four more identical catapults. They compared the distances the balls traveled when launched from a 45° angle to the distances the balls traveled when launched from a 30° angle.
Picture: [Uncaptioned image]
Choices: [Do ping pong balls stop rolling along the ground sooner after being launched from a 30° angle or a 45° angle?, Do ping pong balls travel farther when launched from a 30° angle compared to a 45° angle?, ]
Answer index: 1
Subject: Engineer
Description: Select the Erlenmeyer flask.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 3
@@
Description: iption: Which of the following could Ivan’s test show? Ivan was a landscape architect who was hired to design a new city park. The city council wanted the park to have space for outdoor concerts and to have at least 20% of the park shaded by trees. Ivan thought the concert area should be at least 150 meters from the road so traffic noise didn’t interrupt the music. He developed three possible designs for the park with the concert area in a different location in each design. Then, he tested each design by measuring the distance between the road and the concert area. The passage below describes how the engineering-design process was used to test a solution to a problem. Read the passage. Then answer the question below.
Picture: [Uncaptioned image]
Choices: [if at least 20% of the park would be shaded by trees in each design, which design would have the greatest distance between the concert area and the road, which design would have the least traffic noise in the concert area, ]
Answer index: 1
Subject: Engineer
Description: Select the beaker.
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 3
Subject: Engineer
@@
Description: iption: Identify the question that Zeke’s experiment can best answer. The passage below describes an experiment. Read the passage and then follow the instructions below. Zeke divided 40 unripe bananas evenly among eight paper bags and sealed the bags. He poked 20 small holes in four of the bags and left the other four without holes. He kept the bags at room temperature for three days. Then, Zeke opened the bags and counted the number of brown spots on each banana. He compared the average number of brown spots on bananas from bags with holes to the average number of brown spots on bananas from bags without holes.
Picture: [Uncaptioned image]
Choices: [Do bananas develop more brown spots if they are kept in bags with holes compared to bags without holes?, Do bananas develop more brown spots when they are kept at room temperature compared to in a cold refrigerator?, ]
Answer index: 0
Subject: Engineer
@@
Description: iption: Hint: An independent variable is a variable whose effect you are investigating. A dependent variable is a variable that you measure. Which of the following was an independent variable in this experiment? The passage below describes an experiment. Read the passage and think about the variables that are described. Tyler designed an electric circuit to test how well different types of metal conduct electricity. The circuit included a battery, a light bulb, wires, and clips that could be attached to a sheet of metal. If the metal conducted electricity poorly, the light bulb would appear dim. If the metal conducted electricity well, the light bulb would appear bright. Tyler collected nine equally sized sheets of metal: three sheets of copper, three sheets of iron, and three sheets of aluminum. He used the clips to attach each metal sheet, one sheet at a time, to the circuit. For each sheet, Tyler used a light meter to measure how much light the bulb produced.
Picture: [Uncaptioned image]
Choices: [the amount of light produced by the light bulb, the type of metal sheet used in the circuit, ]
Answer index: 1
Subject: Engineer
@@
Description: iption: Identify the question that Devon’s experiment can best answer. The passage below describes an experiment. Read the passage and then follow the instructions below. Devon poured four ounces of water into each of six glasses. Devon dissolved one tablespoon of salt in each of three glasses, and did not add salt to the other three. Then, Devon placed an egg in one glass and observed if the egg floated. She removed the egg and dried it. She repeated the process with the other five glasses, recording each time if the egg floated. Devon repeated this test with two more eggs and counted the number of times the eggs floated in fresh water compared to salty water.
Picture: [Uncaptioned image]
Choices: [Does the amount of water in a glass affect whether eggs sink or float in the water?, Are eggs more likely to float in fresh water or salty water?, ]
Answer index: 1
Subject: Engineer